OpenStudy (anonymous):

Assume that a procedure yields a binomial distribution with a trial repeated n times. find the probability of x successes given the probability p of success on a single trial. Round to three decimal places. n = 30 x = 12, p = .20

OpenStudy (anonymous):

formula is \[\dbinom{n}{k}p^k(1-p)^{n-k}\]

OpenStudy (anonymous):

in your case \(n=30,k=12,n-k=18,p=.2,1-p=.8\)

OpenStudy (anonymous):

so you need to compute \[\binom{30}{12}(.2)^{12}\times (.8)^{18}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw